OpenStudy (anonymous):

Solve a^2 - 81 = 0 by factoring. Select all possible answers. -9 No solution 9 Infinite solutions 0

OpenStudy (anonymous):

It's 9 and -9. (a+9)(a-9)=0.

OpenStudy (anonymous):

so, the answer is -9 and 9?

OpenStudy (mathstudent55):

You need to factor the difference of two squares. That factors like this: \(a^2 - b^2 = (a + b)(a - b) \) In your case, you have \(a^2 - 81 = 0\) \(a^2 - 9^2 = 0\) \((a + 9 )(a - 9) = 0\) \(a + 9 = 0\) or \(a - 9 = 0\) \(a = -9\) or \(a = 9\)

OpenStudy (mathstudent55):

Yes.

OpenStudy (anonymous):

What about Solve n^2 - 3n = 0 by factoring. Select all possible answers. Infinite solutions No solution Others that aren't listed here. 3 0

OpenStudy (mathstudent55):

Factor out a common term out of the left side. Both terms have n in them.